OpenStudy (lgbasallote):

a 300gal capacity tank contains a solution of 200 gals of water and 50 lbs of salt. a solution containing 3 lbs of salt pergallon is allowed to flow into the tank at teh rate of 4 gpm. the mixture flows from the tank at the rate of 2 gal/min. how many pounds of salt are in the tank at the end of 30 min?

OpenStudy (lgbasallote):

is that important?

OpenStudy (lgbasallote):

well it's not given...it's also not in the formula...

OpenStudy (lgbasallote):

aww

OpenStudy (lgbasallote):

\[\frac{dQ}{dt} = 4(3) - 2[\frac{Q}{200+ (4-2)t}]\] \[\frac{dQ}{dt} = 12 - \frac{2Q}{200 + 2t}\] \[\frac{dQ}{dt} + (\frac{1}{100 + t})Q = 12\] then linear DE thingy i end up with \[Q(100+t) = 6(100+t)^2 + c\]

OpenStudy (asnaseer):

you are told that there are 200 gallons of water with 50lb of salt in it. that mean 50/200 = 1/4 lb of salt per gallon. mixture flows out at a rate of 2 gallons/minute, which means at the end of 30 minutes you would have lost 2*30 = 60 gallons of the original mixture. so we now have 200 - 60 = 140 gallons of 1/4 lb per gallon of salt left from the original mixture. so amount of original salt left = 140 * 1/4 = 35lb mixture flows in at rate of 4 gallons per minute, which means at the end of 30 minutes you would have gained 4*30 = 120 gallons of the new mixture. so we now have 120 gallons of 3lb per gallon of salt added to the original. so amount of new salt added = 120 * 3 = 360lb.

OpenStudy (asnaseer):

so at the end of 30 minutes you have 35lb + 360lb = 395lb of salt in the tank

OpenStudy (lgbasallote):

wait...can you explain this part sir so we now have 200 - 60 = 140 gallons of 1/4 lb per gallon of salt left from the original mixture. so amount of original salt left = 140 * 1/4 = 35lb

OpenStudy (asnaseer):

you agree that mixture flows out at rate of 2 gal/min?

OpenStudy (lgbasallote):

btw it says the answer should be 356.92 lbs...

OpenStudy (asnaseer):

maybe I made a algebra mistake somewhere

OpenStudy (asnaseer):

but see if you can follow my reasoning

OpenStudy (lgbasallote):

there are some parts i dont get :(

OpenStudy (asnaseer):

BTW: I am assuming the mixture flowing in does not /mix/ with the mixture already in the tank. i.e. it just settles at the top.

OpenStudy (lgbasallote):

it mixes

OpenStudy (lgbasallote):

it is being stirred

OpenStudy (asnaseer):

oh - have they given you a formula for how it mixes?

OpenStudy (lgbasallote):

\[\huge \frac{dQ}{dt} = r_1 c_1 - r_2 [\frac{Q}{V_o + (r_1 - r_2)t}]\]

OpenStudy (lgbasallote):

it's the one i was using a while ago

OpenStudy (asnaseer):

what do all the symbols represent?

OpenStudy (lgbasallote):

r1 is the rate of the thingy going in c1 is the lg/gal thingy r2 is the rate of the thingy going out V_o is the volume of water in the tank Q is the amt of salt in the tank

OpenStudy (lgbasallote):

t is time of course

OpenStudy (asnaseer):

ok, so from this (and BTW you really should have stated all this with the question), we get:\[r_1=4gal/min\]\[r_2=2gal/min\]\[c_1=0.25lb/gal\]\[V_0=200gal\]

OpenStudy (asnaseer):

shouldn't there also be a \(c_2\) somewhere?

OpenStudy (asnaseer):

I am assuming \(c_1\) represents the concentration of salt in the original mixture, and \(c_2\) would be the concentration of salt being mixed in?

OpenStudy (asnaseer):

oh - hang on - looks \(Q\) represents the initial concentration of salt in the tank (i.e. at time zero)

OpenStudy (asnaseer):

so \(c_1=3lb/gal\) and \(Q=50lb\) at \(t=0\)

OpenStudy (lgbasallote):

c1 is the salt coming in

OpenStudy (lgbasallote):

the amt of salt coming in

OpenStudy (lgbasallote):

and btw..that formula was a standard for these problems so i thought not to put it anymore

OpenStudy (asnaseer):

yes, so then you end up with:\[\frac{dQ}{dt}=12-\frac{Q}{100+t}\]

OpenStudy (asnaseer):

and we need to integrate this from t=0 to t=30

OpenStudy (lgbasallote):

hmm yep

OpenStudy (lgbasallote):

what i was thinking is to solve for c in that last equation i had \[Q(100 +t) = 6(100+t)^2 + c\] however i dont know how

OpenStudy (asnaseer):

oh - this is a linear DE of the form:\[Q'+(\frac{1}{100+t})Q=12\]

OpenStudy (lgbasallote):

because if i have a value for c then i can just make t = 30 to get Q..idk

OpenStudy (lgbasallote):

yep

OpenStudy (asnaseer):

have you solved these types of DE's? e.g. using an integrating factor?

OpenStudy (lgbasallote):

yup...i solved the DE already... i got \[Q(100 + t) = 6(100+t)^2 + c\]

OpenStudy (asnaseer):

ok - let me check that

OpenStudy (lgbasallote):

okay :)

OpenStudy (asnaseer):

I get a different answer...

OpenStudy (lgbasallote):

but how :O

OpenStudy (asnaseer):

integrating factor = 100 + t

OpenStudy (lgbasallote):

yes

OpenStudy (asnaseer):

which leads to:\[(100+t)Q=\int12(100+t)dt=1200t+6t^2\]giving:\[Q=\frac{6t(200+t)}{100+t}\]

OpenStudy (asnaseer):

and, for t=30, I get Q=318.462

OpenStudy (asnaseer):

oh hang on

OpenStudy (asnaseer):

forgot the constant!

OpenStudy (asnaseer):

\[Q=\frac{6t(200+t)}{100+t}+k\]

OpenStudy (asnaseer):

and we know Q at time t=0

OpenStudy (asnaseer):

Q=50lb at t=0

OpenStudy (lgbasallote):

wait...

OpenStudy (asnaseer):

therefore:\[50=0+k\]therefore k=50

OpenStudy (lgbasallote):

\[\int 12(100 + t)dt\] \[12\int(100+t)dt\] \[12[\frac{(100+t)^2}{2}]\] right?

OpenStudy (asnaseer):

sorry, it should have been:\[Q=\frac{6t(200+t)+k}{100+t}\]

OpenStudy (asnaseer):

no:\[\int(100+t )dt=\int100dt+\int tdt=100t+\frac{t^2}{2}+k\]

OpenStudy (lgbasallote):

hah so my teacher was wrong

OpenStudy (asnaseer):

so we have:\[Q=\frac{6t(200+t)+k}{100+t}\]and Q=50 at t=0, therefore:\[50=\frac{k}{100}\]\[\therefore k=50000\]

OpenStudy (asnaseer):

therefore:\[Q=\frac{6t(200+t)+50000}{100+t}\]

OpenStudy (lgbasallote):

mmhmm

OpenStudy (lgbasallote):

so now t = 30

OpenStudy (asnaseer):

sorry - one too many zeros in there!

OpenStudy (asnaseer):

k=5000

OpenStudy (asnaseer):

\[Q=\frac{6t(200+t)+5000}{100+t}\]therefore, at t=30:\[Q=356.92\]

OpenStudy (lgbasallote):

weirdd...it gave the same answer as when 12(100+t) was used

OpenStudy (lgbasallote):

12(100+t)^2 i mean

OpenStudy (lgbasallote):

i mean 6(100+t)^2

OpenStudy (asnaseer):

really?

OpenStudy (lgbasallote):

yep

OpenStudy (asnaseer):

are you sure - even AFTER computing the constant by using Q=50 at t=0?

OpenStudy (lgbasallote):

well the constant in that one was 55000

OpenStudy (asnaseer):

what was the formula you ended up with?

OpenStudy (lgbasallote):

lol 4th time im gonna type this :)) \[Q(100+t) = 6(100+t)^2 + c\]

OpenStudy (lgbasallote):

so at t = 0 100(50) = 60,000 + c 5000 = 60,000 + c c = -55,000

OpenStudy (lgbasallote):

then at t=30 Q(130) = 6(130)^2 - 55,000

OpenStudy (asnaseer):

that is really wierd

OpenStudy (asnaseer):

let me see if there is a reasonable explanation for that

OpenStudy (lgbasallote):

do you think it's a fluke?

OpenStudy (lgbasallote):

or there really is a calculus reason...

OpenStudy (asnaseer):

both formulas are identical after simplification!

OpenStudy (lgbasallote):

but (100+t)^2/2 = 5,000 + 100 + t^2/2

OpenStudy (lgbasallote):

oh lol i see

OpenStudy (lgbasallote):

thanks for the help sir :D this topic is the only one im having difficulty in our diff eq :/ i get all the other applications except this

OpenStudy (asnaseer):

yw :)

OpenStudy (asnaseer):

I still am amazed that the answers came out the same - because I am sure your integration was not correct :/

OpenStudy (asnaseer):

ah! I see how it happened!

OpenStudy (lgbasallote):

i was following what my teacher did...but i think it's somehow right if i let u = 100+t du = dt so it's int udu

OpenStudy (asnaseer):

look at this: http://www.wolframalpha.com/input/?i=integrate+12%28100%2Bx%29

OpenStudy (asnaseer):

\[\int 12(100 + t)dt=6t^2+1200t+c\]but, you worked it out as:\[\int 12(100 + t)dt=12\frac{(100+t)^2}{2}+c=6(100+t)^2+c=6(100^2+200t+t^2)+c\]\[=6*100^2+1200t+6t^2+c\]\[=6t^2+1200t+c+6*100^2\]

OpenStudy (asnaseer):

so the only difference here is the value of the constant. my constant is just your constant plus 6*100^2

OpenStudy (asnaseer):

your constant came out to be -55000 so -55000 + 6*100^2 = -55000 + 60000 = 5000 = my constant

OpenStudy (asnaseer):

mystery solved!

OpenStudy (lgbasallote):

wow cool

OpenStudy (lgbasallote):

so that proves that in diffeq constants dont matter much

OpenStudy (asnaseer):

yes - and more importantly, your teacher WAS right after all :)